# Enhancing Data Lakes with GraphAr: Efficient Graph Data Management with a Specialized Storage Scheme

### v3

@@ 2.4 Key Challenges Addressed by GraphAr
Challenge 2: Efficient neighbor retrieval. A cornerstone of graph queries is the operation known as neighbor retrieval. This is vital for quick access to adjacent vertices and edges, thus accelerating graph traversal. Existing tabular formats, however, do not natively or efficiently support this crucial operation. This issue is tackled in Section 4.
@@
Figure 4: The metadata and data layout for the example graph in GraphAr.
3 Representing LPGs in GraphAr
This section provides an overview of how GraphAr customizes the representation of LPGs in data lakes. It begins by outlining the goals and non-goals of GraphAr, providing clarity on the rationale behind its design. Next, it explains the strategies employed for data organization and layout, emphasizing the significance of schema metadata and the use of Parquet as the payload format. Lastly, the section describes how GraphAr seamlessly integrates into the data lake ecosystem.
3.1 Goals and Non-Goals
Goals. GraphAr’s primary goal is to provide an efficient storage scheme for LPGs in data lakes, specifically targeting the three main challenges outlined in Section 2.
@@ 3.1 Goals and Non-Goals
In line with the established practices of data warehousing and lake house architectures, both Parquet/ORC and GraphAr adhere to data immutability norms, treating batch-generated data as immutable once created. Higher-level systems, such as graph databases, manage mutation and compaction (e.g., adding, deleting, or updating a vertex) through specialized, non-standardized file and in-memory versioning methods.
3.2 Data Organization and Layout
@@
Schema metadata. A YAML file, as illustrated in Figure 3(a), stores the metadata for a graph. This file specifies important attributes such as file path prefixes, and vertex/edge types. It serves as a nimble yet effective way to capture metadata that is not accommodated by Parquet, while Parquet files include specific details about properties and labels within their internal schemas. The YAML file can optionally include partition sizes, allowing for data to be segmented into multiple physical parquet files, thereby facilitating file-level parallelism.
Vertex table. As depicted in Figure 3(b), each row in the vertex table represents a unique vertex, identified by a 0-indexed internal ID, stored in the <Internal ID> column. Optionally, when partitioning is enabled, for the $i$-th partition, its internal IDs start at partition_size×$\times i$, and within each partition, IDs are sorted in ascending order. Bubbles222“Bubbles” refer to the allowance for some ranges of internal IDs or edge segments not to correspond to any vertices or edges. are allowed at the end of each partition, meaning the actual number of rows can be less than or equal to the partition size.
@@
Property columns (pid and age) are named after their respective properties and hold the corresponding values with specified datatypes. In terms of labels, a set of candidate labels is defined for each vertex type. Then a vertex can have an arbitrary number of labels from the corresponding set. For example, the vertex type Person may have labels to represent ethnicity. For efficient storage and filtering of labels, GraphAr uses a binary representation to maintain each label in an individual column named with angle brackets, e.g., <Asian> and <Enrollee>. Additionally, advanced encoding/decoding techniques are applied, which will be discussed in Section 5.
@@
Edge table. Edges are also organized and stored in Parquet files, similar to vertices. Figure 3(c) showcases the layout of the edge table for type Person-Diagnosed-Disease, where Person and Disease represent the source and destination vertex types, while Diagnosed signifies the classification of the relationships. Each edge is associated with the internal IDs of its source and destination vertices, stored in columns named <src> and <dst>. Edge properties, and optional partitioning, are handled in the same way as vertices.
@@
Optimized access patterns for neighbor retrieval. The layout strategy in GraphAr leverages the columnar storage capabilities of Parquet to facilitate efficient graph traversal. Edges are sorted first by source vertex IDs and then by destination vertice IDs. This sorting strategy optimizes various access patterns. For row-wise access, the layout closely resembles the Coordinate List (COO) format, making it well-suited for edge-centric operations. On the other hand, an auxiliary index table, denoted as <offset>, is introduced to enable more efficient vertex-centric operations. The <offset> table aligns with the partitions in the vertex table, and when applied to the source vertices, facilitates retrieval patterns similar to Compressed Sparse Row (CSR). This separate index table accommodates scenarios where a vertex has multiple types of edges created separately from different data sources, a common occurrence in data lakes. Likewise, a similar approach can be applied to enable Compressed Sparse Column (CSC)-like access. GraphAr allows for efficient retrieval of neighbors in both outgoing and incoming directions, through two sorted tables for the same edge type. CSR, CSC, and COO are widely adopted for representing graphs, by emulating these formats, GraphAr ensures compatibility with existing graph systems.
@@
These layout strategies are complemented by encoding and decoding optimizations, which are detailed in Section 4. Collectively, these strategies serve to enhance both the data management and query capabilities of GraphAr.
3.3 Incorporation with Data Lakes
@@
The design of GraphAr makes it especially well-suited for integration with data lakes, largely due to its reliance on widely adopted standards like Parquet and YAML.
@@
Data transformation and construction. The GraphAr format is essentially a specialized layout of Parquet files accompanied by a YAML metadata file. This enables the use of existing data processing frameworks like Apache Spark, Acero, and Hadoop, which can access various graph systems like Neo4j, TigerGraph and Nebula, or other types of database systems through their respective connectors. These frameworks can also ingest a multitude of data formats including logs, relational tables, JSON, and more. Such flexibility provides users with the ability to construct, transform, and store LPGs in data lakes from a wide array of data sources. To simplify the process of generating files in GraphAr, we have also provided a Spark library specifically designed for this purpose.
@@
4 Efficient Neighbor Retrieval
In this section, we address the critical challenge of efficient neighbor retrieval in graphs. We leverage Parquet’s data pages and introduce page-aligned collections (PAC) for streamlined neighbor identification. Additionally, we harness Parquet’s delta encoding as an efficiency-boosting technique. We then introduce a novel and optimized decoding strategy that leverages BMI and SIMD, using bitmaps as the data structure for collections in PAC. All of these techniques are integrated into GraphAr, resulting in highly efficient neighbor retrieval.
4.1 Workflow and Challenges
@@
For LPG queries, a common operation is to retrieve specific property values of neighboring vertices, given a queried vertex, e.g., obtaining the name values of Disease vertices connected to a particular Person vertex. Assuming the CSR format utilized for storing edge table Person-Diagnosed-Disease, the typical workflow for this operation involves: 1) Using the <offset> index and <dst> column of the edge table to identify and fetch the first relevant page from the target vertex table (a page from the name column in the Disease table). This page contains at least one neighboring vertex pertinent to the query, and may also include other irrelevant vertices; 2) Selectively fetching the property values corresponding to the neighboring vertices within that page. This step is repeated iteratively for each subsequent page containing the targeted neighbors.
This workflow highlights the two primary tasks during neighbor retrieval. The first task is to identify which pages in the vertex table contain the neighboring vertices relevant to the query. The second task is to fetch the relevant property values within each of these pages efficiently. Before discussing our solutions, we formalize the neighbor retrieval operation:
Definition 1 (Page-aligned collections (PAC))
@@
Definition 2 (Neighbor Retrieval)
Given a vertex $v$, the operation of neighbor retrieval returns the PAC $\mathcal{C}$ representing the internal IDs of the neighboring vertices connected to $v$.
@@
Intuition: Each collection $C$ in PAC returned by neighbor retrieval corresponds to a data page, addressing the first task. To save space and avoid unnecessary processing, empty collections are omitted from the PAC. This is based on the sparsity of real-world graphs, which often results in irrelevant pages. For the second task, the internal IDs within each collection enable the retrieval of only the relevant property values through a selection process. Afterwards, the remaining challenges involve efficiently generating PAC $\mathcal{C}$ and optimizing the data structure of each collection for quick value retrieval.
4.2 Delta Encoding
@@
To compute the PAC $\mathcal{C}$, the encoded internal IDs of neighboring vertices need to be loaded, sourced from the edge table. In a data lake scenario, where data can be stored remotely, the loading process can be more time-consuming than processing due to I/O limitations. To address this issue, we investigate the use of delta encoding for data compression, consequently reducing the volume of data that needs to be loaded.
@@
Delta encoding. Previous research, including Gemini [65] and Facebook-Graph [58], has demonstrated that real-world graphs often exhibit both sparsity and locality. This means that while a vertex’s neighbors might be spread across the entire graph, they are more likely to cluster within certain ID ranges. Such patterns arise from various factors, such as the inherent clustering in real-life graph data, where vertices within a cluster are more interconnected, and the methods used for graph data collection (e.g., crawlers or the organic/viral growth patterns of social networks like Facebook or TikTok). Systems [26, 65] have utilized such sparsity and locality to enable efficient partitioning or compression.
@@
In GraphAr, such inherent locality, reinforced by our meticulous dual-key sorting of edges in the carefully designed layout, to enable incremental arrangement of internal IDs for a vertex’s neighbors, serves as the basis for delta encoding, which is highly effective for both the <src> and <dst> columns in the edge table. Delta encoding works by representing the gaps (deltas) between consecutive values instead of storing each value individually. The deltas, which often have small values, can be stored more compactly, requiring fewer bits.
Figure 5: An example of PAC and the bitmap representation.
@@
Figure 6: An example of accelerating decoding via BMI.
Implementations. We utilize Parquet’s built-in support for delta encoding [43], which is implemented based on miniblocks. Each miniblock (with a size of $32$ values) is binary packed using its own bit width, which should be a power of 2 for data alignment purposes. This design allows us to adapt to changes in the data distribution, as the bit width of each miniblock is dynamically adjusted to minimize storage consumption. According to our evaluation across various real-world graphs, as detailed in Section 6.2, the delta encoding technique can reduce the expected loaded data volume (which can be measured by storage consumption) by $58.1\%$ to $81.0\%$ compared to without delta encoding. As a result, it brings an individual speedup of up to 3.6×3.6$\times$3.6 × for neighbor retrieval.
4.3 BMI-based Decoding
Challenges. Whiles delta encoding effectively reduces loading costs, it introduces additional decoding computation. Existing works [45, 43, 61, 50, 51, 52] have explored the use of BMI (Bit Manipulation Instructions) and SIMD (Single Instruction, Multiple Data) to accelerate the data compression, decoding, scanning, or management. However, delta encoding involves data dependencies that make vectorization challenging. The decoding of the $(i+1)$-th neighbor depends on the prior decoding of the $i$-th neighbor.
Another unsolved issue is the choice of implementing collection $C$ in PAC. A pioneering solution proposed in [45] underscores the transformation of indices into a bitmap representation to enable selection pushdown in columnar storages. To adopt this approach to efficiently retrieve the properties of neighbors, which constitutes the second task of neighbor retrieval, we adopt a bitmap representation $B$ for each non-empty collection $C$ in PAC, where $B[i]=1$ indicates the existence of $i$-th element. Figure 5 illustrates a PAC and its bitmap representation, where only $C_{1}$ is non-empty, and $B_{1}$ is the bitmap representation of $C_{1}$. The bitmap representation can be used to facilitate the selection pushdown of vertex properties or labels, e.g., fetching the properties of name in the target vertex table Disease, for the neighbors of a person vertex.
Then, the critical challenge becomes how to efficiently generate the bitmap representation from the delta-encoded neighbor IDs. Existing techniques are not suitable for our context due to the data dependencies involved. However, by taking advantage of the sophisticated instruction sets offered by modern CPUs, we can exploit the functionalities of BMI together with SIMD operations to overcome this challenge, through an innovative decoding strategy.
@@ 4.3 BMI-based Decoding
Acceleration via BMI and SIMD. In practical implementation, the bit-shifting encoding is maintained using a fixed-length datatype, characterized by zero-padding on the left side. In our example, 16 bits are sufficient to accommodate the 4-bit delta values. The bit-shifting encodings of four values are stored in a 64-bit register, allowing for parallel generation and processing using SIMD instructions. Then, the focus shifts to the compaction of these encodings. Fortunately, the Parallel Bit Extract (PEXT) operation, a specialized CPU instruction in BMI, facilitates efficient aggregation of discrete bits from source positions into contiguous bits within the destination, governed by a selector mask, as illustrated in Figure 5(b).
The subsequent challenge is to generate the required mask, achieved by deriving the $i$-th mask $m_{i}$ from the $i$-th bit-shifting encoding $s_{i}$ using the equation $m_{i}=(s_{i}<<1)-1$. The beauty of this operation is its potential for parallel execution, facilitated by direct manipulation of the mask sequence residing within a 64-bit register. Specifically, $M$ can be acquired via the following steps: 1) a bitwise (in our example, 16 bits) shift of each $s_{i}$ to the left by 1 bit, parallelized through SIMD instructions like _mm_slli_epi16; and 2) a bitwise subtraction of 1 from the result of the previous step, which can be accelerated via instructions like _mm_sub_epi16. All these SIMD instructions utilized in our implementations are widely available in modern CPUs, included in SSE2 (which we use) and more recent sets such as AVX2 and AVX-512.
In general, vectorization demonstrates greater efficiency when the bit width is smaller, as it allows for more significant parallelism. Our extensive evaluations have confirmed that the BMI-based decoding approach outperforms the default decoding approach in Parquet when the bit width is within 4 bits, with performance improvements ranging from $3.2\%$ to $110\%$. Therefore, we utilize this BMI-based approach for miniblocks with a bit width of 1 to 4 bits, while resorting to the default delta decoding of Parquet for larger bit widths. The combination of data layout, delta encoding, and this adaptive decoding strategy results in an advanced topology management paradigm, enabling efficient neighbor retrieval.
5 Optimized Label Filtering
Labels serve as a representation of the classification or characteristics of vertices in a graph. Filtering vertices by labels is a fundamental syntax in graph query languages, as it allows querying specific subsets of vertices. However, existing approaches [8, 15] of fitting graphs into tabular data often treat labels as regular properties, encapsulating them within a string or list, as seen in Figure 3. These approaches overlook the inherent differences between labels and properties, leading to inefficiencies when performing label filtering, due to the need for decoding string representations and string matching.

### v4

@@ 2.4. Key Challenges Addressed by GraphAr
Challenge 2: Efficient neighbor retrieval. A fundamental aspect of graph queries is the operation known as neighbor retrieval. This is vital for quick access to adjacent vertices and edges, thus accelerating graph traversal. Existing tabular formats, however, do not natively or efficiently support this crucial operation. This issue is tackled in Section 4.
@@
Figure 4. The metadata and data layout for the example graph in GraphAr.
3. Representing LPGs in GraphAr
This section provides an overview of how GraphAr customizes the representation of LPGs in data lakes. It begins by outlining its goals and non-goals, providing clarity on the rationale behind its design. Next, it explains the strategies employed for data organization and layout. Lastly, the section describes how GraphAr seamlessly integrates into the data lake ecosystem.
3.1. Goals and Non-Goals
Goals. GraphAr’s primary goal is to provide an efficient storage and management scheme for LPGs in data lakes, specifically targeting the three main challenges outlined in Section 2.
@@ 3.1. Goals and Non-Goals
Also, GraphAr is not a graph computing engine; rather, it can be non-intrusively integrated with graph processing systems, either serving as the archival format or acting as a data source.
In line with the established practices of data warehousing and lake house architectures, both Parquet/ORC and GraphAr adhere to data immutability norms, treating batch-generated data as immutable once created. Higher-level systems such as graph databases manage mutation (e.g., adding, deleting, or updating vertices) through specialized, non-standardized file and in-memory versioning methods.
3.2. Data Organization and Layout
@@
Schema metadata. A YAML file (Figure 4(a)) stores the metadata for a graph. It specifies important attributes such as file path prefixes and vertex/edge types. This file serves as a nimble yet effective way to capture metadata that is not accommodated by Parquet, while Parquet files include specific details about properties and labels within their internal schemas. It can optionally include partition sizes, allowing for data to be segmented into multiple physical Parquet files, thereby enabling parallelism at the file level.
[…]
@@
Property columns (pid and age) are named after their respective properties and hold the corresponding values with specified datatypes. In terms of labels, a set of candidate labels is defined for each vertex type. Then a vertex can have an arbitrary number of labels from the corresponding set. For example, the vertex type Person may have labels to represent ethnicity. For efficient storage and filtering of labels, GraphAr uses a binary representation to maintain each label in an individual column named with angle brackets, e.g., ¡Asian¿ and ¡Enrollee¿. Additionally, advanced encoding/decoding techniques are applied, which will be discussed in Section 5.
@@
Edge table. Edges are also organized and stored in Parquet files, similar to vertices. Figure 4(c) showcases the layout of the edge table for type Person-Diagnosed-Disease, where Person and Disease represent the source and destination vertex types, while Diagnosed signifies the classification of the relationships. Each edge is associated with the internal IDs of its source and destination vertices, stored in columns named ¡src¿ and ¡dst¿. Edge properties, and optional partitioning, are handled in the same way as vertices.
@@
Optimized access patterns for neighbor retrieval. The layout strategy in GraphAr leverages the columnar storage capabilities of Parquet to facilitate efficient graph traversal. Edges are sorted first by source vertex IDs and then by destination vertex IDs. This sorting strategy optimizes various access patterns. For row-wise access, the layout closely resembles the Coordinate List (COO) format, making it well-suited for edge-centric operations. On the other hand, an auxiliary index table, denoted as ¡offset¿, is introduced to enable more efficient vertex-centric operations. The ¡offset¿ table aligns with the partitions in the vertex table, and when applied to the source vertices, facilitates retrieval patterns similar to Compressed Sparse Row (CSR). Likewise, a similar approach can be applied to enable Compressed Sparse Column (CSC)-like access. GraphAr allows for efficient bidirectional neighbor retrieval through two sorted tables for the same edge type. CSR, CSC, and COO are widely adopted for representing graphs, thus GraphAr ensures compatibility with existing graph-related systems.
@@
These layout strategies are complemented by encoding and decoding optimizations (Section 4). Collectively, these strategies enhance both the data management and query capabilities of GraphAr.
3.3. Incorporation with Data Lakes
@@
The design of GraphAr makes it especially well-suited for integration with data lakes, largely due to its reliance on widely adopted standards such as Parquet and YAML.
@@
Data transformation and construction. The GraphAr format is essentially a specialized layout of Parquet files accompanied by a YAML metadata file. This enables the use of existing data processing frameworks like Apache Spark, Acero, and Hadoop, which can access various graph systems like Neo4j, TigerGraph and Nebula, or other types of database systems through their respective connectors. These frameworks can also ingest a multitude of data formats including logs, relational tables, JSON, and more. Such flexibility provides users with the ability to construct, transform, and store LPGs in data lakes from a wide array of data sources.
@@
4. Efficient Neighbor Retrieval
In this section, we address the critical challenge of efficient neighbor retrieval in graphs. We leverage Parquet’s data pages and introduce page-aligned collections (PAC) for streamlined neighbor identification. Additionally, we utilize Parquet’s delta encoding to enhance performance, and introduce an innovative decoding strategy that leverages BMI and SIMD. All of these techniques are integrated into GraphAr, resulting in highly efficient neighbor retrieval.
4.1. Workflow of Neighbor Retrieval
@@
For LPG queries, a common operation is to retrieve specific property values of neighboring vertices, given a queried vertex, e.g., obtaining the name values of Disease vertices connected to a particular Person vertex. Assuming the CSR format utilized for storing edge table Person-Diagnosed-Disease, the typical workflow for this operation involves: 1) Using the ¡offset¿ index and ¡dst¿ column of the edge table to identify and fetch the first relevant page from the target vertex table (a page from the name column in the Disease table). This page contains at least one neighboring vertex pertinent to the query, and may also include other irrelevant vertices; 2) Selectively fetching the property values corresponding to the neighboring vertices within that page. This step is repeated iteratively for each subsequent page containing the targeted neighbors.
This workflow highlights the two primary steps during neighbor retrieval. The first is to identify which pages in the vertex table contain the neighboring vertices relevant to the query. The second is to fetch the relevant property values within each of these pages efficiently. Firstly, we formalize the neighbor retrieval operation:
Definition 0 (Neighbor Retrieval).
Given a vertex $v$, the operation of neighbor retrieval returns a data structure $\mathcal{C}$ representing the internal IDs of the neighboring vertices connected to $v$.
Requirement of $\mathcal{C}$. The data structure $\mathcal{C}$ should facilitate the retrieval of the relevant property values of neighboring vertices, while minimizing space and processing overhead. Accordingly, we introduce the concept of page-aligned collections (PAC).
4.2. Page-Aligned Collections
Definition 0 (Page-aligned collections (PAC)).
@@ 4.2. Page-Aligned Collections
Intuition. Each collection $C$ in PAC returned by neighbor retrieval corresponds to a data page of the target vertex table. To save space and avoid unnecessary processing, empty collections are omitted. This is based on the sparsity and locality of real-world graphs, which often results in irrelevant pages. Subsequently, the internal IDs within each collection enable the retrieval of only the relevant property values through a selection process. The remaining challenges then involve 1) optimizing each collection’s representation for quick value retrieval and 2) efficiently generating the PAC $\mathcal{C}$.
A pioneering solution (Li et al., 2023) underscores the transformation of indices into a bitmap representation to enable selection pushdown in columnar storages. To adopt this approach to efficiently retrieve the properties of neighbors, which addresses the first challenge, we adopt a bitmap representation $B$ for each non-empty collection $C$ in PAC, where $B[i]=1$ indicates the existence of $i$-th element.
Figure 5(a) illustrates the neighbor retrieval of a source vertex to obtain a PAC, and Figure 5(b) demonstrates the usage of PAC to get the properties of its neighbors. For this illustrated example, only $C_{0}$ is non-empty, and $B_{0}$ is the bitmap representation of $C_{0}$. The bitmap representation can be used to facilitate the selection pushdown of vertex properties or labels, e.g., fetching the properties of name in the target vertex table Disease, for the neighbors of Person $\text{vertex}_{0}$. For the second challenge of efficiently generating PAC (i.e., Figure 5(a)), we utilize Parquet’s delta encoding and a novel decoding strategy, which we detail in the following.
4.3. Delta Encoding
@@
To compute the PAC $\mathcal{C}$, the encoded internal IDs of neighboring vertices need to be loaded, sourced from the edge table. In a data lake scenario, where data can be stored remotely, the loading process can be more time-consuming than processing due to I/O limitations. To address this issue, we investigate the use of delta encoding for data compression, consequently reducing data load volume.
@@
Delta encoding. Previous research (Zhu et al., 2016; Ugander et al., 2011) has demonstrated that real-world graphs often exhibit both sparsity and locality. This means that while a vertex’s neighbors might be spread across the entire graph, they are more likely to cluster within certain ID ranges. Such patterns arise from various factors, such as the inherent clustering in real-life graphs, where vertices within a cluster are more interconnected, and the methods used for data collection (e.g., crawlers or the organic/viral growth patterns of social networks like Facebook or TikTok). Systems (Boldi and Vigna, 2004; Zhu et al., 2016) have utilized such sparsity and locality to enable efficient partitioning or compression.
@@
In GraphAr, such inherent locality, reinforced by our meticulous dual-key sorting of edges in the carefully designed layout, to enable incremental arrangement of internal IDs for a vertex’s neighbors, serves as the basis for delta encoding, which is highly effective for both the ¡src¿ and the ¡dst¿ columns in the edge table. Delta encoding works by storing the deltas between consecutive values rather than each value separately. The deltas, which often have small values, can be stored more compactly, requiring fewer bits.
(a) Neighbor retrieval of a source vertex to obtain PAC.
(b) Using the PAC to get the properties of its neighbors.
Figure 5. An example of PAC and its usage.
@@
Figure 6. An example of accelerating decoding via BMI.
Implementations. We utilize Parquet’s built-in support for delta encoding (Lemire and Boytsov, 2013), which is implemented based on miniblocks. Each miniblock (with a size of $32$ values) is binary packed using its own bit width, which should be a power of 2 for data alignment purposes. This design allows us to adapt to changes in the data distribution, as the bit width of each miniblock is dynamically adjusted to minimize storage consumption. According to our evaluation across various real-world graphs, as detailed in Section 6.2, the delta encoding technique can reduce the expected loaded data volume by $58.1\%$ to $81.0\%$ compared to without delta encoding. As a result, it brings an individual speedup of 2.7×2.7$\times$2.7 × for neighbor retrieval.
4.4. BMI-based Decoding
Challenges. Whiles delta encoding effectively reduces loading costs, it introduces additional decoding computation. Some existing works (Li et al., 2023; Lemire and Boytsov, 2013; Willhalm et al., 2009; Polychroniou et al., 2015; Polychroniou and Ross, 2019, 2020) have explored the use of BMI (Bit Manipulation Instructions) and SIMD (Single Instruction, Multiple Data) to accelerate the data compression, decoding, scanning, or management. However, delta encoding involves data dependencies that make vectorization challenging. The decoding of the $(i+1)$-th neighbor depends on the prior decoding of the $i$-th neighbor.
In our context, the critical challenge is to generate the bitmap representation of PAC efficiently from the delta-encoded neighbor IDs, which are stored in the ¡dst¿ column of the edge table (as shown in Figure 6(a)). Existing techniques are not suitable for our context due to the data dependencies involved. However, by taking advantage of the sophisticated instruction sets offered by modern CPUs, we can exploit the functionalities of BMI together with SIMD operations to overcome this challenge, through an innovative decoding strategy.
@@ 4.4. BMI-based Decoding
Acceleration via BMI and SIMD. In practical implementation, the bit-shifting encoding is maintained using a fixed-length datatype, characterized by zero-padding on the left side. In our example, 16 bits are sufficient to accommodate the 4-bit deltas. The bit-shifting encodings of 4 values are stored in a 64-bit register, allowing for parallel generation and processing using SIMD instructions. Then, the focus shifts to the compaction of these encodings. Fortunately, the Parallel Bit Extract (PEXT) operation, a specialized CPU instruction in BMI, facilitates efficient aggregation of discrete bits from source positions into contiguous bits within the destination, governed by a selector mask. This process is illustrated in Figure 6(b).
The subsequent challenge is to generate the required mask, achieved by deriving the $i$-th mask $m_{i}$ from the $i$-th bit-shifting encoding $s_{i}$ using the equation $m_{i}=(s_{i}<<1)-1$. The advantage of this operation is its potential for convenient parallel execution, facilitated by direct manipulation of the mask sequence residing within a 64-bit register. Specifically, $M$ can be acquired via the following steps: 1) a bitwise (in our example, 16 bits) shift of each $s_{i}$ to the left by 1 bit, parallelized through SIMD instructions like _mm_slli_epi16; and 2) a bitwise subtraction of 1 from the result of the previous step, which can be accelerated via instructions like _mm_sub_epi16. All these SIMD instructions utilized in our implementations are widely available in modern CPUs, included in SSE2 (which we use) and more recent sets such as AVX2 and AVX-512.
In general, vectorization demonstrates greater efficiency when the bit width is smaller, as it allows for more significant parallelism. Our extensive evaluations have confirmed that the BMI-based decoding approach outperforms the default decoding approach in Parquet when the bit width is within 4 bits, with performance improvements ranging from $3.3\%$ to $110\%$. Therefore, we utilize this BMI-based approach for miniblocks with a bit width of 1 to 4 bits, while resorting to the default delta decoding of Parquet for larger bit widths. The combination of data layout, delta encoding, and this adaptive decoding strategy results in an advanced topology management paradigm, enabling efficient neighbor retrieval.
5. Optimized Label Filtering
Labels serve as a representation of the classification or characteristics of vertices in a graph. Filtering vertices by labels is a fundamental syntax in graph query languages, as it allows querying specific subsets of vertices. Existing approaches (neo, 2024a, d) of fitting graphs into tabular data often treat labels as regular properties, encapsulating them within a string or list, as seen in Figure 3(a). This approach overlooks the inherent differences between labels and properties, leading to inefficient label filtering, due to the need for decoding string representations and conducting string matching.